If there is 3 cups of sugar for every 2 cup of vinegar how much sugar for 1 cup of vinegar.
Answer by Theo(13342)

ratio of sugar to vinegar is 3/2.

you want to keep that ratio the same.

your equation is:

3/2 = x/1

cross multiply to get 2x = 3

divide both sides of the equation by 2 to get x = 3/2.

you would need 3/2 cup of sugar for 1 cup of vinegar.

the ratio of 3/2 is the same as the ratio of (3/2) / 1